# Dynamical Lensing Tomography of Black Hole Ringdown

### v2

@@ I Introduction
One of the foundational tests of General Relativity concerns the deflection of light by massive bodies Will (2014); Lemos et al. (2019); Bartelmann and Schneider (2001); Bartelmann (2010): the path of light rays is “bent” by mass-energy, which distorts the spacetime fabric. In the last decade, it has become clear that the universe is flooded with traveling spacetime distortions, gravitational waves (GWs) traveling at the speed of light. Gravitational-wave astronomy provides a wealth of information on compact objects, particularly on the behavior of gravity in highly dynamical, high-curvature and redshift setups.
An intriguing prospect concerns the detectability of light bending by GWs, which would add an extra information channel on strong gravity. This possibility was raised and studied by different authors in the weak field regime Labeyrie (1993); Fakir (1994); Durrer (1994); Kaiser and Jaffe (1997); Damour and Esposito-Farese (1998). Our purpose here is to study light deflection in the strong-field, highly dynamical regime. In particular, we will study the appearance of a black hole (BH) which is approaching a quiescent state via relaxation in its characteristic modes of oscillation, called quasinormal modes (QNMs). The relaxation stage is also referred to as ringdown stage, and is a generic late-time description of the coalescence of compact objects Berti et al. (2009); Baibhav et al. (2023). Our results could describe stars on the background of a binary BH merger, or the appearance of a circumbinary accretion disk surrounding two merging compact objects.
@@ II Ray Tracing towards a Relaxing Black Hole
A relaxing BH is well described by general relativistic first-order perturbation theory, where spacetime is expressed by small fluctuations away from the final, stationary state. Uniqueness results in General Relativity suggest that the final state, to a good approximation, is a Kerr BH Chrusciel et al. (2012); Cardoso and Gualtieri (2016); Cardoso et al. (2022). We thus take our relaxing BH spacetime, of mass $M$ and angular momentum $aM$ to be given by
@@
where $g_{\mu\nu}^{\mathrm{(0)}}$ is the Kerr metric (we will use Boyer-Lindquist $(t,r,\theta,\phi)$ coordinates throughout Bardeen et al. (1972)). The term $h_{\mu\nu}$ denotes first-order perturbations.
Instead of working directly with the metric perturbations $h_{\mu\nu}$, we consider instead curvature perturbations Teukolsky (1973, 1972). These can be separated and decoupled and all the relevant information about radiative degrees of freedom is encapsulated in a master variable
@@
The gravitational quantity of interest is $\Psi_{4}^{(1)}=\psi_{-2}(t,r,\theta,\phi)(r-ia\cos\theta)^{-4}$, the first-order perturbation of one of the Weyl scalars, which is directly related to the amplitude of GWs far from the source. Notation and conventions are shown in Supplemental Material A 1. The function ${}_{s}S(\theta)$ is a spin-weighted spheroidal harmonic, while the radial Teukolsky function ${}_{s}R(r)$ obeys a homogeneous second-order differential equation. We fix its normalization by insisting on the large-distance behavior. Then, ${}_{s}Z$ controls the amplitude of the GW, by matching the ringdown stage to some physically generated process. The characteristic modes are found by imposing proper boundary conditions on ${}_{s}R(r)$ , which selects only a discrete set of frequencies $\omega$, the ringdown frequencies Berti et al. (2009). All the functions and quantities carry a dependence on angular numbers $\ell,m$ and a possible overtone number $n$ Teukolsky (1973, 1972); Berti et al. (2009). For simplicity, we do not include these indices in the subscripts. Although a true GW signal comprises a superposition of various modes, our discussion focuses on the dominant $\ell=m=2$ mode. Additionally, we exclusively address the $n=0$ mode, which lasts longer compared to the more rapidly decaying higher overtones. For QNMs characterized by frequency $\omega$ and labeled by $(\ell,m,n)$, there exists a corresponding “mirror” mode with frequency $-\bar{\omega}$ and $(\ell,-m,n)$. In the final spin frame, these paired modes primarily represent GW emissions propagating in opposite directions along the spin axis—one mode predominantly radiating towards the northern hemisphere and the other towards the southern hemisphere Li et al. (2022); Berti et al. (2006); Cook and Zalutskiy (2014); Krivan et al. (1997). The observed waveform typically manifests as a superposition of these two distinct sets of modes. However, one of the paired modes may be less prominent due to weaker excitation Berti et al. (2006); Hamilton et al. (2023). In the following analysis, we primarily focus on the $m=2$ mode with $\mathrm{Re}(\omega)>0$, which predominantly emits in the southern hemisphere, as more general cases exhibit similar features.
@@
Once ${}_{-2}R(r)$ is known, we reconstruct the metric fluctuation $h_{\mu\nu}$ following well-known prescriptions Chrzanowski (1975); Cohen and Kegeles (1974); Kegeles and Cohen (1979), summarized in Supplemental Material A 2. The QNMs diverge as $h_{\mu\nu}\sim e^{\mathrm{Im}(\omega)(t-r_{*})}$ when $t-r_{*}\to-\infty$, where $\mathrm{Im}(\omega)$ represents the imaginary part of the ringdown frequency $\omega$ and $r_{*}$ is the tortoise coordinate Berti et al. (2009). To mitigate the divergence, we truncate the perturbation metric $h_{\mu\nu}$ by applying a Heaviside-like filter function $H(t-r_{*})$:
$H(t-r_{*})=\frac{1}{2}\left(\tanh{\frac{t-r_{*}}{q}}+1\right)\,,$ (3)
to obtain $H(t-r_{*})\,h_{\mu\nu}$. Here, $0<q<-2/\mathrm{Im}(\omega)$, where $q$ controls the intermediate region width of the filter. Although various values of $q$ were tested, the modifications had only a minor impact on the outcome, especially regarding the late-time frequency behavior. In the following, we employ a relatively sharp truncation in the filter function $H(t-r_{*})$, with $-q\mathrm{Im}(\omega)=0.1$.
@@
To characterize the observable effects, it is essential to define the observer frame. In an unperturbed Kerr BH scenario, a zero angular momentum observer (ZAMO) frame is typically used Frolov and Novikov (1998). However, in a dynamic spacetime, the ZAMO tetrads are neither orthogonal nor normalized. To rectify this, we apply a Gram-Schmidt orthogonalization process to establish a normalized, orthogonal tetrad basis derived from the ZAMO tetrads, as detailed in Supplemental Material B 1. Using this newly defined observer frame, we initiate geodesics towards the BH employing backward ray tracing. Here, each pixel on the observer plane, represented by coordinates $(\tilde{X},\tilde{Y})$ Bardeen (1973); Gralla et al. (2018); Gralla and Lupsasca (2020), corresponds to photon momentum in this frame. The photon’s 44 44-momentum in the dynamic spacetime $p^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ is then constructed from this orthogonal tetrad basis, similar to the procedure established in Kerr spacetime Cunha et al. (2016). For convenience, we also introduce polar coordinates for the observer plane centered on the direction directly facing the BH in Kerr spacetime, denoted as $(\tilde{\rho},\tilde{\varphi})$, where $\tilde{\rho}$ is the impact parameter and $\tilde{\varphi}$ is the angle.
@@ III Lensing Tomography
Figure 1: Top: Illustration of photon trajectories in a relaxing Kerr BH spacetime with spin $a=0.7M$, viewed from a face-on observer with $\theta_{0}\approx 0$. The metric perturbation is modeled using the $n=0$ and $\ell=m=2$ mode, with magnitude normalized to $|{}_{-2}Z|=0.24$. The initial impact parameters are set at $\tilde{\rho}=5.041M$, slightly exceeding the critical curve $\tilde{\rho}_{c}=5.04M$ in Kerr spacetime. The black line represents unperturbed geodesics in Kerr spacetime, an arrow indicating the direction from the observer to the BH to depict the backward ray tracing method used. Different colors represent geodesics launched at various observation times $t_{0}-r_{*}(r_{0})$, detailed in the bottom panel. The dashed line marks a trajectory terminating on the BH, whereas the others extend towards infinity. The gray plane indicates the BH equatorial plane. Bottom: Growth of geodesic deviation $\Delta_{\rm KS}$ representing the distance between unperturbed and deflected geodesics at the same backward propagating time $t_{0}-t-r_{*}(r_{0})$, shown in Cartesian KS coordinates.
@@
We explore geodesics in a relaxing Kerr spacetime, particularly focusing on the merger events of supermassive binary BHs (SMBHBs) with equal mass ratios. The remnant BH from each mass ratio merger events typically characterized by an angular momentum parameter $a\sim 0.7\,M$ Buonanno et al. (2007), a value consistent with numerous observations Nitz et al. (2019); Abbott et al. (2019) and considered as our fiducial value. The GW signal from these mergers is dominated by the $\ell=|m|=2$ mode Berti et al. (2007, 2009, 2006); Maggiore (2018), with the $n=0$ mode persisting the longest. This fundamental mode is characterized by $M\omega=0.5326-0.08079i$ Berti et al. (2009); Stein (2019).
@@
In the vicinity of the BH, the metric perturbation amplitude can reach up to $M^{2}\Psi_{4}^{(1)}\sim 0.1$ Berti et al. (2007). The total radiated energy during the ringdown is limited to approximately $3\%$ of the remnant BH’s mass Berti et al. (2007), setting an upper limit on the magnitude of $|{}_{-2}Z|$ for the $\ell=m=2$ mode at about 0.240.24 0.240.24. We adopt this upper bound as the fiducial value for our analyses, our results can be trivially re-scaled for generic amplitudes.
@@
We first consider a face-on observer with $\theta_{o}\approx 0$. The top panel of Fig. 1 illustrates backward photon trajectories, all launched from the same direction, using identical backward impact parameters relative to the observer positioned at $r_{0}=10^{9}M$. These trajectories are depicted using the Cartesian Kerr–Schild (KS) coordinate system $(x,y,z)$. The black line represents geodesics in an unperturbed Kerr spacetime, while various colors indicate different emission times $t_{o}-r_{*}(r_{o})$. The bottom panel displays the growth of deviations between unperturbed and deflected geodesics, defined as $\Delta_{\rm KS}\equiv(\Delta x^{2}+\Delta y^{2}+\Delta z^{2})^{1/2}$, representing the distance between the two trajectories at the same backward propagating time $t_{0}-t-r_{*}(r_{o})$, colors corresponding to trajectory lines in the top panel.
@@
The evolution of geodesic deviations, launched at different observer times, initially shows a weak increase after departing from the observer. Geodesics launched later exhibit deviations exponentially smaller than those launched earlier, reflecting the exponential decay of GW amplitude near the BH during the ringdown phase. Closer to the BH, all deviations begin to exhibit exponential growth due to the instability of the radial potential in the light ring region Podurets (1965); Ames and Thorne (1968); Luminet (1979); Campbell and Matzner (1973); Cunningham and Bardeen (1972); Falcke et al. (2000); Chandrasekhar (1983); Cardoso et al. (2009); Johnson et al. (2020); Gralla and Lupsasca (2020); Cardoso et al. (2021); Chen et al. (2023). Here, the interplay between light ring instability and GW waves induces addiontial oscillation features in the exponential growth. The geodesics that depart from the light ring region begin to develop deviations that increase linearly with time.
@@
Figure 2: Variations of the deflected angles in the azimuthal ($\Delta\phi$) and polar ($\Delta\theta$) components for geodesics terminating at infinity, plotted as a function of observation time $t_{0}-r_{*}(r_{o})$. The BH, observer, and ringdown configuration follow the setup described in Fig. 1. Different colors represent geodesics launched from various angles $\tilde{\varphi}$ on the polar coordinate of the observer plane, with impact parameters fixed at $\tilde{\rho}=5.4M$. Variations for cases separated by $\pi$ in $\tilde{\varphi}$ appear identical, reflecting the $m=2$ nature of the BH ringdown.
@@
Figure 2 displays the deflected angles in polar $(\Delta\theta)$ and azimuthal $(\Delta\phi)$ components as a function of observer frame time, $t_{0}-r_{*}(r_{o})$, for trajectories initially launched from various polar coordinate angles $\tilde{\varphi}$ on the observer plane. The impact parameters are fixed at $\tilde{\rho}=5.4M$. Due to ringdown metric perturbations, axial symmetry is not preserved even in face-on observations. Notably, both $\Delta\phi$ and $\Delta\theta$ show identical variations for values of $\tilde{\varphi}$ separated by $\pi$. This pattern is precisely attributable to the $m=2$ mode of the metric perturbations, akin to those generated by vector or tensor superradiant clouds Chen et al. (2023), demonstrating that the deflected angles can effectively dissect the polarization pattern of the GWs around the BH.
Next, we adjust the impact parameters $\Delta\tilde{\rho}\equiv\tilde{\rho}-\tilde{\rho}_{c}$ relative to the critical curve $\tilde{\rho}_{c}\approx 5.04M$, as defined in Kerr spacetime. In Fig. 3, we illustrate variations of the deflected angle $\tilde{\delta}$. Due to the implementation of the filter, deflection begins around $t_{0}-r_{*}(r_{o})\approx 0$. For impact parameters close to the critical curve, specifically in cases where $\Delta\tilde{\rho}\leq 1.5M$, the early time evolution of the deflected angle exhibits a pattern of exponential decay and oscillation. The timescales of this decay and the oscillation period align with the inverses of the imaginary and real parts of $\omega$, respectively, suggesting that these observables can effectively capture key characteristics of the BH ringdown.
Furthermore, when comparing the overall amplitude of $\tilde{\delta}$, there is a noticeable exponential increase dependent on $\Delta\tilde{\rho}$. This enhancement is again attributed to the instability of the BH light ring region, which causes an exponential growth in the perturbed deviation during propagation around the BH Podurets (1965); Ames and Thorne (1968); Luminet (1979); Campbell and Matzner (1973); Cunningham and Bardeen (1972); Falcke et al. (2000); Chandrasekhar (1983); Cardoso et al. (2009); Johnson et al. (2020); Gralla and Lupsasca (2020); Chen et al. (2023). As $\tilde{\rho}$ approaches the critical curve, this instability intensifies. The instability is quantified by the square root of the second derivative of the radial potential Johnson et al. (2020), which increases linearly as $\Delta\tilde{\rho}$ decreases, thereby leading to an exponential rise in the deflected angle.

### v3

@@ I Introduction
One of the foundational tests of General Relativity concerns the deflection of light by massive bodies [1, 2, 3, 4]: the path of light rays is “bent” by mass-energy, which distorts the spacetime fabric. In the last decade, it has become clear that the universe is flooded with traveling spacetime distortions, gravitational waves (GWs) traveling at the speed of light. Gravitational-wave astronomy provides a wealth of information on compact objects, particularly on the behavior of gravity in highly dynamical, high-curvature and redshift setups.
An intriguing prospect concerns the detectability of light bending by GWs, which would add an extra information channel on strong gravity. This possibility was raised and studied by different authors in the weak field regime [5, 6, 7, 8, 9]. Our purpose here is to study light deflection near compact objects in highly dynamic spacetimes. In particular, we will study the appearance of a black hole (BH) which is approaching a quiescent state via relaxation in its characteristic modes of oscillation, called quasinormal modes (QNMs). The relaxation stage is also referred to as ringdown stage, and is a generic late-time description of the coalescence of compact objects [10, 11]. Our results could describe stars on the background of a binary BH merger, or the appearance of a circumbinary accretion disk surrounding two merging compact objects.
@@ II Ray Tracing towards a Relaxing Black Hole
A relaxing BH is well described by general relativistic first-order perturbation theory, where spacetime is expressed by small fluctuations away from the final, stationary state. Uniqueness results in General Relativity suggest that the final state, to a good approximation, is a Kerr BH [12, 13, 14]. We thus take our relaxing BH spacetime, of mass $M$ and angular momentum $aM$ to be given by
@@
where $g_{\mu\nu}^{\mathrm{(0)}}$ is the Kerr metric (we will use Boyer-Lindquist $(t,r,\theta,\phi)$ coordinates throughout [15]). The term $h_{\mu\nu}$ denotes first-order perturbations.
Instead of working directly with the metric perturbations $h_{\mu\nu}$, we consider instead curvature perturbations [16, 17]. These can be separated and decoupled and all the relevant information about radiative degrees of freedom is encapsulated in a master variable
@@
The gravitational quantity of interest is $\Psi_{4}^{(1)}=\psi_{-2}(t,r,\theta,\phi)(r-ia\cos\theta)^{-4}$, the first-order perturbation of one of the Weyl scalars, which is directly related to the amplitude of GWs far from the source. Notation and conventions are shown in Supplemental Material A.1 [18]. The function ${}_{s}S(\theta)$ is a spin-weighted spheroidal harmonic, while the radial Teukolsky function ${}_{s}R(r)$ obeys a homogeneous second-order differential equation. We fix its normalization by calculating energy fluxes at infinity, and comparing to numerical relativity results. Then, ${}_{s}Z$ controls the amplitude of the GW, by matching the ringdown stage to some physically generated process. The characteristic modes are found by imposing proper boundary conditions on ${}_{s}R(r)$ , which selects only a discrete set of frequencies $\omega$, the ringdown frequencies [10]. All the functions and quantities carry a dependence on angular numbers $\ell,m$ and a possible overtone number $n$ [16, 17, 10]. For simplicity, we do not include these indices in the subscripts. A GW signal comprises a superposition of various modes, but the $\ell=|m|=2$ dominates (so far it is the only one detected beyond any doubt). Additionally, we exclusively address the $n=0$ mode, which lasts longer compared to the more rapidly decaying higher overtones. For similar reasons we neglect nonlinearities, which are of much smaller amplitude even for equal-mass mergers, and of shorter duration [19, 20]. For QNMs characterized by frequency $\omega$ and labeled by $(\ell,m,n)$, there exists a corresponding “mirror” mode with frequency $-\bar{\omega}$ and $(\ell,-m,n)$. In the final spin frame, these paired modes primarily represent GW emissions propagating in opposite directions along the spin axis [21, 22, 23, 24]. The observed waveform typically manifests as a superposition of these two distinct sets of modes. However, one of the paired modes may be less prominent due to weaker excitation [22, 25]. In the following analysis, we primarily focus on the $m=2$ mode with $\mathrm{Re}(\omega)>0$, which predominantly emits toward the southern hemisphere, as more general cases exhibit similar features, as discussed in Supplemental Material C [18].
@@
Once ${}_{-2}R(r)$ is known, we reconstruct the metric fluctuation $h_{\mu\nu}$ following well-known prescriptions [26, 27, 28], summarized in Supplemental Material A.2 [18]. We initiate the ringdown at $t=r_{*}$, where $r_{*}$ denotes the tortoise radial coordinate, by applying a Heaviside-like filter to the perturbation, $h_{\mu\nu}\to H(t-r_{*})\,h_{\mu\nu}$, along the light cone, with
$H(t-r_{*})=\frac{1}{2}\left(\tanh{\frac{t-r_{*}}{\chi}}+1\right)\,.$ ) . (3)
In the following, we employ a relatively sharp truncation in the filter function $H(t-r_{*})$, with $-\chi\mathrm{Im}(\omega)=0.1$.
@@
Figure 1: Schematic diagram of the coordinate system. Left: Illustration of a light ray originating from the observer (green point) and traced backward (black arrow) toward the BH, terminating at the celestial sphere plane at infinity (blue plane). The black and red lines represent the geodesics in unperturbed Kerr spacetime and those deflected by metric perturbations, respectively. The grand arc difference is denoted by $\tilde{\delta}$, with angular variations in the azimuthal and polar directions labeled as $\Delta\phi$ and $\Delta\theta$. Right: Diagram of the observer plane with polar coordinates $(\tilde{\rho},\tilde{\varphi})$. The critical curve $\tilde{\rho}_{c}(\tilde{\varphi})$ is shown as a black contour, and the relative impact parameter is defined as $\Delta\tilde{\rho}\equiv\tilde{\rho}-\tilde{\rho}_{c}$.
To characterize the observable effects, it is essential to define the observer frame. In an unperturbed Kerr BH scenario, a zero angular momentum observer (ZAMO) frame is typically used [29]. However, in a dynamic spacetime, the ZAMO tetrads are neither orthogonal nor normalized. To rectify this, we apply a Gram-Schmidt orthogonalization process to establish a normalized, orthogonal tetrad basis derived from the ZAMO tetrads, as detailed in Supplemental Material B.1 [18]. Using this newly defined observer frame, we initiate geodesics towards the BH employing backward ray tracing. Here, each pixel on the observer plane, represented by coordinates $(\tilde{X},\tilde{Y})$ [30, 31, 32], corresponds to photon momentum in this frame. The photon’s 44 44-momentum in the dynamic spacetime $p^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ is then constructed from this orthogonal tetrad basis, similar to the procedure established in Kerr spacetime [33]. For convenience, we also introduce polar coordinates for the observer plane centered on the direction directly facing the BH in Kerr spacetime, denoted as $(\tilde{\rho},\tilde{\varphi})$, where $\tilde{\rho}$ is the impact parameter and $\tilde{\varphi}$ is the angle.
@@ III Lensing Tomography
Figure 2: Growth of geodesic deviation $\Delta_{\rm KS}$, representing the distance between unperturbed and deflected geodesics at the same backward-propagating time $\tilde{t}_{o}-t$, in a relaxing Kerr BH spacetime with spin $a=0.7M$, viewed from a face-on observer with $\theta_{0}\approx\pi$, is shown in Cartesian KS coordinates. Different colors represent geodesics launched at various observation times $\tilde{t}_{o}$, as detailed in the accompanying inset panel. The dashed line marks a trajectory terminating on the BH, whereas the others extend towards infinity. The metric perturbation is modeled using the $n=0$ and $\ell=m=2$ mode, with the magnitude normalized to $|{}_{-2}Z|=0.24$. The initial impact parameters are set at $\tilde{\rho}=5.041M$, slightly exceeding the critical curve $\tilde{\rho}_{c}=5.04M$ in Kerr spacetime. Inset: Illustration of the corresponding photon trajectories. The black line represents unperturbed geodesics in Kerr spacetime, with an arrow indicating the direction from the observer to the BH to depict the backward ray-tracing method used. The gray plane indicates the BH equator.
@@
We explore geodesics in a relaxing Kerr spacetime, particularly focusing on the merger events of supermassive binary BHs (SMBHBs) with equal mass ratios. The remnant BH from each mass ratio merger events typically characterized by an angular momentum parameter $a\sim 0.7\,M$ [34], a value consistent with numerous observations [35, 36] and considered as our fiducial value. The GW signal from these mergers is dominated by the $\ell=|m|=2$ mode [37, 10, 22, 38], with the $n=0$ mode persisting the longest. This fundamental mode is characterized by $M\omega=0.5326-0.08079i$ [10, 39].
@@
In the vicinity of the BH, the metric perturbation amplitude can reach up to $M^{2}\Psi_{4}^{(1)}\sim 0.1$ [37]. The total radiated energy during the ringdown, $E_{\mathrm{tot}}$, is limited to approximately $3\%$ of the remnant BH’s mass [37], setting an upper limit on the magnitude of $|{}_{-2}Z|$ for the $\ell=m=2$ mode at about 0.240.24 0.240.24. We adopt this upper bound as the fiducial value for our analyses, our results can be trivially re-scaled for generic amplitudes.
@@
We first consider a face-on observer with $\theta_{o}\approx\pi$. Figure 2 illustrates backward photon trajectories, all launched from the same direction, using identical backward impact parameters relative to the observer positioned at $r_{0}=10^{9}M$. These trajectories are depicted using the Cartesian Kerr–Schild (KS) coordinate system $(x,y,z)$. The black line represents a geodesic in an unperturbed Kerr spacetime, while the various colors indicate different observation times $\tilde{t}_{o}\equiv t_{o}-r_{*}(r_{o})$, where $\tilde{t}_{o}=0$ is approximately the time when the GW ringdown initially reaches the observer. The bottom panel displays the growth of deviations between unperturbed and deflected geodesics, defined as $\Delta_{\rm KS}\equiv(\Delta x^{2}+\Delta y^{2}+\Delta z^{2})^{1/2}$, representing the distance between the two trajectories at the same backward propagating time $\tilde{t}_{o}-t$, colors corresponding to trajectory lines in the inset panel.
@@
The evolution of geodesic deviations, launched at different observer times, initially shows a weak increase after departing from the observer. Geodesics launched later exhibit deviations exponentially smaller than those launched earlier, reflecting the exponential decay of GW amplitude near the BH during the ringdown phase. Closer to the BH, all deviations begin to exhibit exponential growth due to the instability of the radial potential in the light ring region [40, 41, 42, 43, 44, 45, 46, 47, 48, 32, 49, 50]. Here, the interplay between light ring instability and GW induces additional oscillation features in the exponential growth. The geodesics that depart from the light ring region begin to develop deviations that increase linearly with time.
@@
Figure 3: Evolutions of the deflected angles in the azimuthal ($\Delta\phi$) and polar ($\Delta\theta$) components for geodesics terminating at infinity, plotted as a function of observation time $\tilde{t}_{o}$. The configuration of the BH, observer, and ringdown is consistent with the setup described in Fig. 2. Different colors illustrate geodesics launched from various angles on the polar coordinate of the observer’s plane $\tilde{\varphi}$, all with fixed impact parameters of $\tilde{\rho}=6M$. The symmetrical patterns observed in cases separated by $\pi$ in $\tilde{\varphi}$ highlight the $m=2$ symmetry of the BH ringdown.
@@
Figure 3 displays the deflected angles in polar $(\Delta\theta)$ and azimuthal $(\Delta\phi)$ components as a function of observer frame time, $\tilde{t}_{o}$, for trajectories initially launched from various polar coordinate angles $\tilde{\varphi}$ on the observer plane. The impact parameters are fixed at $\tilde{\rho}=6M$. Due to ringdown metric perturbations, axial symmetry is not preserved even in face-on observations. Notably, both $\Delta\phi$ and $\Delta\theta$ show identical variations for values of $\tilde{\varphi}$ separated by $\pi$. This pattern is precisely attributable to the $m=2$ mode of the metric perturbations, akin to those generated by vector or tensor superradiant clouds [50], demonstrating that the deflected angles can effectively dissect the polarization pattern of the GWs around the BH.
In the top panel of Fig. 4, we depict the evolution of the deflection angle $\tilde{\delta}$ for different relative impact parameters $\Delta\tilde{\rho}\equiv\tilde{\rho}-\tilde{\rho}_{c}$, where $\tilde{\rho}_{c}\approx 5.04M$ is defined as the critical curve in Kerr spacetime (for $a=0.7M$). The early time evolutions of $\tilde{\delta}$ consistently exhibit a pattern of exponential decay and oscillation that aligns with the time-dependence of the gravitational waveform shown in the bottom panel. This alignment suggests that these observables effectively capture key characteristics of the BH ringdown. We fit the envelope of the $\Delta\tilde{\rho}/M=0.5$ case using the ringdown decay rate, specifically, $\tilde{\delta}\sim A_{\tilde{\delta}}\exp(\mathrm{Im}(\omega)\,\tilde{t}_{o})$, as shown in the dot-dashed line. This demonstrates a clear overlap in the early stages. The amplitude of the deflection angle $A_{\tilde{\delta}}$ is well described at large $\Delta\tilde{\rho}$ by
$A_{\tilde{\delta}}\sim\frac{1.2\sqrt{E_{\mathrm{tot}}/M}}{\Delta\tilde{\rho}/M}\,,$ (5)
with $E_{\mathrm{tot}}/M=0.03$ in our fiducial case.